OpenStudy (anonymous):

How do i set up this? find the domain of square root of 2x-10 by finding the intervals on the real number line?

OpenStudy (anonymous):

the domain is the set of all Xs that are valid since the equation is \[\sqrt{2x-10}\] and you can not have a - square root you need to find all values of X that satisfy the equation: \[2x -10 \ge 0 \] when you solve for X on that equation you get \[x \ge 5\] so the domain of this function is \[x \ge 5\]

OpenStudy (anonymous):

Thanks!!